OpenStudy (anonymous):

the value pi/4 is a solution for the equation 3 sqrt2 sec theta +7 =1 true or false?

OpenStudy (alexwee123):

try plugging it in

OpenStudy (anonymous):

what do I do to plug it in?

OpenStudy (anonymous):

\[3\sqrt{2}\sec{\theta}+7=1\] put \(\frac{\pi}{4}\) where \(\theta\) is: \[3\sqrt{2}\sec\left(\frac{\pi}{4}\right)+7=1\] Now, are both sides equal?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

Correct.